# Stationary solution to Stochastically Forced Euler-Poisson Equations in Bounded Domain: Part 1. 3-D Insulating Boundary

### v2

@@ 2.1. Local existence
Similarly as in [21, 28], we first linearize the system, then we use Banach’s fixed point theorem to get the local existence by the a priori energy estimates.
@@
$\mathcal{D}(\hat{\sigma})\operatorname{d}_{t}\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,1}+\tilde{\mathcal{A}}^{2}(\hat{\mathbf{w}})\mathbf{w}_{,2}+\tilde{\mathcal{A}}^{3}(\hat{\mathbf{w}})\mathbf{w}_{,3}+\tilde{\mathcal{B}}\mathbf{w}\right)\operatorname{d}t$ (2.74)
$=$ ~$-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\left(\hat{\sigma},{\bf u}\right)\operatorname{d}t+\tilde{\mathcal{L}}_{\hat{\phi}}\left(\hat{\sigma},\hat{\phi}\right)\operatorname{d}t+\tilde{f}(\hat{\mathbf{w}}).$
where $\hat{\mathbf{w}}=\left[\begin{array}[]{l}\hat{\sigma}\\ \hat{{\bf u}}\end{array}\right]$ is given, $\hat{\sigma}\in C\left([0,T];H^{3}\left(U\right)\right)$, $\hat{{\bf u}}\in C\left([0,T];H^{3}\left(U\right)\right)$, and we denote $M=\sup\limits_{t\in[0,T]}\left\|\hat{\sigma},\hat{{\bf u}}\right\|_{3}$.
Step 1: Uniform bound.
@@
$\int_{U}\operatorname{d}_{t}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x=\int_{U}\frac{1}{2}\operatorname{d}_{t}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\operatorname{d}x+\int_{U}\mathcal{D}\mathbf{w}\cdot\operatorname{d}_{t}\mathbf{w}\operatorname{d}x+\int_{U}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t.$ (2.75)
@@
$\mathcal{D}\operatorname{d}_{t}\mathbf{w}\cdot\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}t$ (2.76)
@@
$\int_{U}\mathcal{D}\mathbf{w}\cdot\operatorname{d}_{t}\mathbf{w}\operatorname{d}x$
@@
$+\int_{U}\nabla h\left(\hat{\sigma}\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+\int_{U}\mathcal{D}\mathbb{F}\operatorname{d}_{t}W\cdot\mathbf{w}\operatorname{d}x.$
@@
$\int_{U}\operatorname{d}_{t}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x$
$=$ $\int_{U}\frac{1}{2}\operatorname{d}_{t}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\operatorname{d}x-\int_{U}\left(-\frac{1}{2}\left(\mathbf{w}\tilde{\mathcal{A}}^{1}_{1}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{2}_{2}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{3}_{3}\mathbf{w}\right)+\tilde{\mathcal{B}}\left|\mathbf{w}\right|^{2}\right)\operatorname{d}x\operatorname{d}t$ (2.80)
$+\int_{U}\left(-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\cdot\mathbf{w}+\tilde{\mathcal{L}}_{\hat{\phi}}\cdot\mathbf{w}\right)\operatorname{d}x\operatorname{d}t+\int_{U}\nabla h\left(\hat{\sigma}\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+\int_{U}\mathcal{D}\mathbb{F}\operatorname{d}_{t}W\cdot\mathbf{w}\operatorname{d}x$

### v3

@@ 2.1. Local existence
In this subsection, the main estimates for the stochastic forces are taking the assumption of (1.20) for instance. Similar to the approach in [37, 45], we first linearize the system and then we use Banach’s fixed point theorem to get the local existence by the a priori energy estimates.
@@
$\mathcal{D}(\hat{\sigma})\operatorname{d}\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,1}+\tilde{\mathcal{A}}^{2}(\hat{\mathbf{w}})\mathbf{w}_{,2}+\tilde{\mathcal{A}}^{3}(\hat{\mathbf{w}})\mathbf{w}_{,3}+\tilde{\mathcal{B}}\mathbf{w}\right)\operatorname{d}t$ (2.74)
$=$ ~$-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\left(\hat{\sigma},{\bf u}\right)\operatorname{d}t+\tilde{\mathcal{L}}_{\hat{\phi}}\left(\hat{\sigma},\hat{\phi}\right)\operatorname{d}t+\tilde{f}(\hat{\mathbf{w}}),$
where $\hat{\mathbf{w}}=\left[\begin{array}[]{l}\hat{\sigma}\\ \hat{{\bf u}}\end{array}\right]$ is given, $\hat{\sigma}\in C\left([0,T];H^{3}\left(U\right)\right)$, $\hat{{\bf u}}\in C\left([0,T];H^{3}\left(U\right)\right)$. We denote $M=\sup\limits_{t\in[0,T]}\left\|\hat{\sigma},\hat{{\bf u}}\right\|_{3}$.
Step 2: Estimates for the uniform upper bound.
@@
$\int_{U}\operatorname{d}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x=\int_{U}\frac{1}{2}\operatorname{d}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\operatorname{d}x+\int_{U}\mathcal{D}\mathbf{w}\cdot\operatorname{d}\mathbf{w}\operatorname{d}x+\int_{U}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t.$ (2.75)
@@
$\mathcal{D}\operatorname{d}\mathbf{w}\cdot\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}t$ (2.76)
@@
$\int_{U}\mathcal{D}\mathbf{w}\cdot\operatorname{d}\mathbf{w}\operatorname{d}x$
@@
$+\int_{U}\nabla h\left(\hat{\sigma}\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+\int_{U}\mathcal{D}\mathbb{F}\operatorname{d}W\cdot\mathbf{w}\operatorname{d}x.$
@@
$\int_{U}\operatorname{d}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x$
$=$ $\int_{U}\frac{1}{2}\operatorname{d}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\operatorname{d}x-\int_{U}\left(-\frac{1}{2}\left(\mathbf{w}\tilde{\mathcal{A}}^{1}_{1}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{2}_{2}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{3}_{3}\mathbf{w}\right)+\tilde{\mathcal{B}}\left|\mathbf{w}\right|^{2}\right)\operatorname{d}x\operatorname{d}t$ (2.80)
$+\int_{U}\left(-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\cdot\mathbf{w}+\tilde{\mathcal{L}}_{\hat{\phi}}\cdot\mathbf{w}\right)\operatorname{d}x\operatorname{d}t+\int_{U}\nabla h\left(\hat{\sigma}\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+\int_{U}\mathcal{D}\mathbb{F}\operatorname{d}W\cdot\mathbf{w}\operatorname{d}x$